the distance between a 60-kg boy and a 90-kg girl is 2.3 m.what is the magnitude of the gravitational force excerts on the other.universal constant=6.67 × 10-11 N m2/ kg2


1
Expert's answer
2021-03-10T17:16:38-0500
"F=G\frac{mM}{r^2}\\F=(6.67\cdot10^{-11})\frac{60\cdot 90}{2.3^2}=6.8\cdot10^{-8}\ N"